OpenStudy (anonymous):

Find the x intercepts of the following: y = x^2 + 6x − 16

OpenStudy (anonymous):

x= -8

OpenStudy (anonymous):

(x-2)(x+8)=0

OpenStudy (anonymous):

rather x= -8,2

OpenStudy (anonymous):

Thank you very much~ :D

hero (hero):

0 = x^2 + 8x -2x -16 0 = x(x+8)-2(x+8) 0 = (x+8)(x-2)